The branching of a blood vessel is shown in Fig. B.8. Blood at a pressure of 18.5 kPa flows in the main vessel at 4.5 L/min. Determine the blood pressure in each branch, assuming that blood vessels behave as rigid tubes, the flow is frictionless, and that the vessel lies in a horizontal plane. You may approximate blood to have a SG of 1.06.

02 dz=3mm (2) di=1cm Qi-4.5L/min P=18.5 kPa (1) d=5mm Q:=2L/min (3) Q3 Fig. B.8
Answer #1

dq=3 mm -5 Is = 7.5015xlo X105n2 Solution :- jog dizlem Q=4.54m2 Pi=10.5 kpa 03=2L/min dz=5mm 01 = Og taz Qq = 4.5-2=2,5 L/miNOW Using Bernoulles theorm V2 Pa tz, + + Zg Yw ag Yw zg 2 Pi V2 + Z1=22 ag qx9,81 7 P2 = P, + v, - Ve Yw = 18.5+ 0.955²5.896